# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 2.4 The Iwahori Hecke algebra of $G$
In this subsection we summarize some results of [IM65].
Fix a maximal torus $T$ of $G$ and an Iwahori subgroup $I$ of $G$ such that $T^{0}\subset I$.
@@
[…]
@@
    if $s_{1},s_{2}\in\tilde{\Delta}$ $s_{1}\neq s_{2}$ and $(s_{1}s_{2})^{m}=1$ then $T_{s_{1}}T_{s_{2}}T_{s_{1}}...=T_{s_{2}}T_{s_{1}}T_{s_{2}}...$ where each product contains $m$ elements.
@@ 2.5 Involutions and symmetric subgroups
We call $X=G/H$ a $p$-adic symmetric space. A $p$-adic symmetric space is an $l$ space (See 1.1 in [BZ76] for a definition).
@@
An apartment $\mathcal{A}$ is called $\sigma$ stable if $\sigma(\mathcal{A})=\mathcal{A}$, notice that $\sigma$ stable maximal tori correspond to $\sigma$ stable apartments under the correspondence mention in 2.11.
@@
Let $L\subset\mathcal{B}_{G}$ be a $\sigma$ stable set with non empty interior which is contained in some apartment. Then $L$ is contained in a $\sigma$ stable apartment.
@@
Let $\mathcal{C}$ be a chamber, $\sigma(\mathcal{C})$ is another chamber, let $L=\mathcal{C}\cup\sigma(\mathcal{C})$, by theorem 2.12, $L$ is contained in some apartment. $L$ has a non empty interior as $\mathcal{C}$ has a non empty interior.
@@
Consider the action of $H$ on the building $\mathcal{B}_{G}$ of $G$. Translating to the language of the building we can deduce from Section 6 in [HW93] the following:
@@ 3 Iwahori orbits on symmetric spaces
In this section we give a description of the set of $I$ orbits on $X$ and define an action of $W_{aff}$ on them.
By theorem 2.15, the set $I\backslash G$ is in bijection with the set of $\Omega$ colored chambers in the affine building of $G$. The orbits of $I$ on $X$ are the same as the orbits of $H$ on $I\backslash G$ or on the $\Omega$ colored chambers of $\mathcal{B}_{G}$.
In order to describe the $H$ orbits on $I\backslash G$ we will need the following proposition about the $\sigma$ stable apartments.
@@
Let $\mathcal{B}_{G}$ be the enlarged affine building of $G$, let $\mathcal{C}$ be a chamber of $\mathcal{B}_{G}$, then there exists a $\sigma$ stable apartment $\mathcal{A}$ that contains $\mathcal{C}$ and such apartment is unique up to the action of an element of $H$ that fixes $\mathcal{C}$ pointwise.
Proof.
The existence is Corollary 2.34 in the present paper. It is not stated in [DS11] but as explained in our preliminaries, it is a direct consequence of a main result of that paper.
We now prove the uniqueness part.
Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be two $\sigma$ stable apartments that contain $\mathcal{C}$. By theorem 2.12 item 3, there is $g\in G$ such that $g\mathcal{A}_{1}=\mathcal{A}_{2}$ and $g$ fixes $S=\mathcal{A}_{1}\cap\mathcal{A}_{2}$ pointwise.
Let $T_{1}$ be the maximal torus that corresponds to $\mathcal{A}_{1}$ and $T_{2}$ be the maximal torus that corresponds to $\mathcal{A}_{2}$.
@@
By item 4 of theorem 2.13 and the Iwahori decomposition (Theorem 2.19) the group of elements that fixes $S$ pointwise is $P_{S}=U_{S}T_{1}^{0}$.
$T_{1}^{0}$ acts trivially on $\mathcal{A}_{1}$ so we may assume that $g\in U_{S}$, $\sigma$ maps $U_{S}$ to itself as it maps $S$ to itself and being an algebraic map sends unipotent elements to unipotent elements. So $\sigma(g)\in U_{S}$ hence $g^{-1}\sigma(g)\in U_{S}$, on the other hand $g^{-1}\sigma(g)$ maps $\mathcal{A}_{1}$ to $\mathcal{A}_{1}$ so $g^{-1}\sigma(g)\in N_{G}(T_{1})$.
@@
Elements of $N_{G}(T_{1})\cap U_{S}$ act trivially on $\mathcal{A}_{1}$. Therefore, $N_{G}(T_{1})\cap U_{S}\subset T_{1}^{0}\cap U_{S}$. Let $T^{-}$ be the sub-torus of $T_{1}$ on which $\sigma$ acts as inverse, i.e. $T^{-}=\{t\in T_{1}|\sigma(t)=t^{-1}\}$. Clearly, $\sigma$ acts on $g^{-1}\sigma(g)$ as inverse. Therefore, $g^{-1}\sigma(g)\in(T^{-})^{0}\cap U_{S}$. $T$ splits over $F$ and so does $T^{-}$. There is $r\in\mathbb{N}$ such that $T^{-}\cong(F^{\times})^{r}$ and $(T^{-})^{0}\cong(\mathcal{O}^{\times})^{r}$. Choose an isomorphism $\phi:(T^{-})^{0}\rightarrow(\mathcal{O}^{\times})^{r}$.
@@
We have $\phi(U_{S}\cap(T^{-})^{0})\subset(\mathcal{O}^{1})^{r}$, for a proof see for example Lemma 7.3.16 in [KP23], where a much more general result is proven. By Hensel’s Lemma, any element of $(\mathcal{O}^{1})^{r}$ has a square root inside $(\mathcal{O}^{1})^{r}$. Thus ,we can find $t\in(\mathcal{O}^{1})^{r}$ such that $\phi(g^{-1}\sigma(g))=t^{2}$, let $g^{\prime}=\phi^{-1}(t)\in(T^{-})^{0}$. We have $g^{-1}\sigma(g)=(g^{\prime})^{2}=g^{\prime}\sigma(g^{\prime})^{-1}$. Another way to write this is $gg^{\prime}=\sigma(gg^{\prime})$. Thus $gg^{\prime}\in H$. We know that $g^{\prime}\in T^{0}$, so $g^{\prime}$ acts trivially on $\mathcal{A}_{1}$. Thus, $gg^{\prime}\mathcal{A}_{1}=g\mathcal{A}_{1}=\mathcal{A}_{2}$ and $gg^{\prime}$ acts on $\mathcal{C}$ as $g$ does which means it fixes it pointwise and we are done.
@@ 3.1 I orbits on $X$
To describe the $H$ orbits on $I\backslash G$ we remind that $I\backslash G$ can be identified with the set of $\Omega$ colored chambers in the affine building $\mathcal{B}_{G}$. We denote $\Omega$ colored chambers by $C_{\Omega}(G)=\{(\mathcal{C},o)|\mathcal{C}\subset\mathcal{B}_{G},o\in\Omega\}$ $G$ acts by $g(\mathcal{C},o)=(g\mathcal{C},\omega(g)o)$.
Let $\mathbf{A}_{\sigma}$ be the set of $\sigma$ stable apartments, $H$ acts on $\mathbf{A}_{\sigma}$ and by Theorem 2.35 this action has finitely many orbits. Let $\mathcal{A}$ be a representative of an orbit and let $C_{\mathcal{A}}$ be the set of all $\Omega$ colored chambers contained in an apartment that is in the $H$ orbit of $\mathcal{A}$. Clearly $C_{\mathcal{A}}$ does not depend on the chosen representative and is stable under the action of $H$. By Proposition 3.1, the sets of the form $C_{\mathcal{A}}$ are disjoint.
So it is enough to describe the $H$ orbits on $C_{\mathcal{A}}$.
@@
The map that sends $W_{H,aff}(T)w$ to the $H$ orbit of $(w\mathcal{C},\omega(w))$ is a bijection between
$W_{H,Aff}(T)\backslash W_{Aff}(T)$ and the $H$ orbits on $C_{\mathcal{A}}$.
@@
It is clear that the map is well defined and surjective, we only have to check that the map is injective. Assume we have $w_{1},w_{2}\in W_{aff}(T)$ such that there is $h\in H$ with $hw_{1}\mathcal{C}=w_{2}\mathcal{C}$ and $\omega(hw_{1})=\omega(w_{2})$, we need to show that $w_{1}w_{2}^{-1}\in HT^{0}$.
Let $h\in H$ such that $hw_{1}\mathcal{C}=w_{2}\mathcal{C}$ and $\omega(hw_{1})=\omega(w_{2})$, we have $hw_{1}w_{2}^{-1}w_{2}\mathcal{C}=w_{2}\mathcal{C}$. By Proposition 3.1 applied to the chamber $w_{2}\mathcal{C}$ we know that there is $h^{\prime}\in H$ that fixes $w_{2}\mathcal{C}$ pointwise such that $hw_{1}w_{2}^{-1}\mathcal{A}=h^{\prime}\mathcal{A}$.
Therefore $h^{\prime-1}hw_{1}w_{2}^{-1}$ stabilizes $\mathcal{A}$ and fixes the chamber $w_{2}\mathcal{C}$ pointwise (pointwise, because $\omega(h^{\prime})=\omega(hw_{1}w_{2}^{-1})=1)$. Thus $h^{\prime-1}hw_{1}w_{2}^{-1}\in T^{0}$. This finishes the proof. ∎
For any $\sigma$ stable torus $T$ denote $W^{H}_{aff}(T)=W_{H,aff}(T)\backslash W_{aff}(T)$.
@@
Let $I$ be the group that fixes a chamber that is contained in $\mathcal{A}_{0}$, for any $i\neq 0$ choose $g_{i}$ such that $g_{i}\mathcal{A}_{0}=\mathcal{A}_{i}$ and set $g_{0}=1$, let $T$ be the torus that corresponds to $\mathcal{A}_{0}$.
From the previous proposition we deduce:
@@
3.2 Action of $W_{aff}$ on $H\backslash G/I$
Theorem 3.3 allows us to define an action of the affine Weyl group of $G$ on the $I$ orbits on $X$. The group acts separately on each $W^{H}_{aff}(T)$ and on each one it acts via the isomorphism between the abstract group $W_{aff}$ and $W_{aff}(T)$, and the action of $W_{aff}(T)$ on its quotient.
We denote by $w\times x$ the resulting action of $w\in W_{aff}$ on $x\in I\backslash G/H$. The action doesn’t depend on the choices of the representatives $g_{i}$.
The choice of the isomorphism between $W_{aff}$ and $W_{aff}(T)$ is the same as the choice of a chamber in proposition 3.2 and doesn’t change the action of the group.
@@ 3.2 Action of $W_{aff}$ on $H\backslash G/I$
Notice that by the construction of the action it follows that for $x\in I\backslash G/H$ and $w\in W_{aff}$ we can find a $\Omega$ colored chamber $(\mathcal{C}_{x},o_{x})\in C_{\Omega}(G)$ whose $H$ orbit is $x$ and another $\Omega$ colored chamber $(\mathcal{C}_{w\times x},o_{w\times x})\in C_{\Omega}(G)$ whose $H$ orbits is $w\times x$ such that $\mathcal{C}_{x}$ and $\mathcal{C}_{w\times x}$ are contained in a single $\sigma$ stable apartment.

### v3

@@ 2.4 The Iwahori Hecke algebra of $G$
In this subsection we summarize some results from [18].
Fix a maximal split torus $T$ of $G$ and an Iwahori subgroup $I$ of $G$ such that $T^{0}\subset I$.
@@
The multiplication operation is given by convolution with respect to a Haar measure $m$ on $G$, $f*g(y)=\int f(yx^{-1})g(x)dm(x)$.
@@
    if $s_{1},s_{2}\in\tilde{\Delta}$ such that $s_{1}\neq s_{2}$ and $(s_{1}s_{2})^{m}=1$ then $T_{s_{1}}T_{s_{2}}T_{s_{1}}...=T_{s_{2}}T_{s_{1}}T_{s_{2}}$... where each product contains $m$ elements.
@@ 2.5 Involutions and symmetric subgroups
[…]
@@
An apartment $\mathcal{A}$ is called $\sigma$ stable if $\sigma(\mathcal{A})=\mathcal{A}$, notice that $\sigma$ stable maximal tori correspond to $\sigma$ stable apartments under the correspondence mentioned in 2.11.
@@
Let $L\subset\mathcal{B}_{G}$ be a $\sigma$ stable set with a non empty interior which is contained in some apartment. Then $L$ is contained in a $\sigma$ stable apartment.
@@
Let $\mathcal{C}$ be a chamber, $\sigma(\mathcal{C})$ is another chamber, let $L=\mathcal{C}\cup\sigma(\mathcal{C})$. By Theorem 2.12, $L$ is contained in some apartment. The set $L$ has a non empty interior as $\mathcal{C}$ has a non empty interior and $L$ is $\sigma$ stable.
@@
Consider the action of $H$ on the building $\mathcal{B}_{G}$ of $G$. Translating to the language of the building we can deduce from Proposition 6.10 and Corollary 6.16 of [15] the following:
@@
Definition 2.37.
Let $S(X)^{I}$ be the space of $I$ invariant locally constant compactly supported functions on $X$. The algebra $H(G,I)$ acts from the right on $S(X)^{I}$. The action is by convolution with respect to a Haar measure $m$.
For $f\in S(X)^{I}$ and $h\in H(G,I)$ we define $f*h(x)=\int f(xg^{-1})h(g)dm(g)$.
Proposition 2.38.
Definition 2.37 gives a well defined right action of $H(G,I)$ on $S(X)^{I}$.
Proof.
We need to check that for $f\in S(X)^{I}$ and $h_{1},h_{2}\in H(G,I)$ we have $(f*h_{1})*h_{2}=f*(h_{1}*h_{2})$.
We have $f*h_{1}(x)=\int f(xg_{1}^{-1})h_{1}(g_{1})dm(g_{1})$ and so
$(f*h_{1}(x))*h_{2}=\int\int f(xg^{-1}_{2}g_{1}^{-1})h_{1}(g_{1})dm(g_{1})h_{2}(g_{2})dm(g_{2})$.
Denote $g_{3}=g_{1}g_{2}$, we have $(f*h_{1})*h_{2}(x)=\int\int f(xg^{-1}_{3})h_{1}(g_{3}g_{2}^{-1})dm(g_{3})h_{2}(g_{2})dm(g_{2})=\int f(xg^{-1}_{3})\int h_{1}(g_{3}g_{2}^{-1})h_{2}(g_{2})dm(g_{2})dm(g_{3})=\int f(xg^{-1}_{3})(h_{1}*h_{2})(g_{3})dm(g_{3})=f*(h_{1}*h_{2})(x).$
@@ 3 Iwahori orbits on symmetric spaces
In this section, we give a description of the set of $I$ orbits on $X$ and define an action of $W_{aff}$ on them.
By Theorem 2.15, the set $G/I$ is in bijection with the set of $\Omega$ colored chambers in the affine building of $G$. The orbits of $I$ on $X$ are the same as the orbits of $H$ on $G/I$ or on the set of $\Omega$ colored chambers in $\mathcal{B}_{G}$.
In order to describe the $H$ orbits on $G/I$ we will need the following proposition about the $\sigma$ stable apartments.
@@
Let $\mathcal{B}_{G}$ be the enlarged affine building of $G$, let $\mathcal{C}$ be a chamber of $\mathcal{B}_{G}$, then there exists a $\sigma$ stable apartment $\mathcal{A}$ that contains $\mathcal{C}$, and such an apartment is unique up to the action of an element of $H$ that fixes $\mathcal{C}$ pointwise.
Remark 3.2.
We can formulate the above proposition without the language of the building. The equivalent formulation is as follows. Let $I^{\prime}$ be any Iwahori subgroup of $G$, there exists a $\sigma$ stable maximal split torus $T^{\prime}$ such that the maximal compact subgroup $T^{\prime 0}$ of $T^{\prime}$ is contained in $I^{\prime}$, i.e. $T^{\prime 0}\subset I^{\prime}$. Such a torus is unique up to conjugation by $H\cap I^{\prime}$.
This is an analogue of Lemma 2.4 of [15].
Proof of Proposition 3.1.
The existence follows from Corollary 2.35 in the present paper. Although this is not stated in [12], as explained in the preliminaries, it is a direct consequence of a main result of that paper.
We now prove uniqueness.
Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be two $\sigma$ stable apartments that contain $\mathcal{C}$. By item 3 of Theorem 2.12, there exists $g\in G$ such that $g\mathcal{A}_{1}=\mathcal{A}_{2}$ and $g$ fixes $S=\mathcal{A}_{1}\cap\mathcal{A}_{2}$ pointwise.
Let $T_{1}$ be the maximal torus corresponding to $\mathcal{A}_{1}$ and let $T_{2}$ be the maximal torus corresponding to $\mathcal{A}_{2}$.
@@
By item 4 of Theorem 2.13 and the Iwahori decomposition (Theorem 2.20) the group of elements that fix $S$ pointwise is $P_{S}=U_{S}T_{1}^{0}$.
$T_{1}^{0}$ acts trivially on $\mathcal{A}_{1}$ so we may assume that $g\in U_{S}$. The involution $\sigma$ maps $U_{S}$ to itself as it maps $S$ to itself and being an algebraic map sends unipotent elements to unipotent elements. So $\sigma(g)\in U_{S}$, hence $g^{-1}\sigma(g)\in U_{S}$. Furthermore, $g^{-1}\sigma(g)$ maps $\mathcal{A}_{1}$ to $\mathcal{A}_{1}$ so $g^{-1}\sigma(g)\in N_{G}(T_{1})$.
@@
Elements of $N_{G}(T_{1})\cap U_{S}$ act trivially on $\mathcal{A}_{1}$. Therefore, $N_{G}(T_{1})\cap U_{S}\subset T_{1}^{0}\cap U_{S}$.
The torus $T_{1}$ is split, so there is some $r\in\mathbb{N}$ and an isomorphism $\phi:T_{1}\rightarrow(F^{\times})^{r}$. Using $\phi$ we can define an involution $\phi\circ\sigma\circ\phi^{-1}$ on $(F^{\times})^{r}$. We abuse the notation and denote this involution also by $\sigma$.
The map $\phi$ also induces an isomorphism between $T_{1}^{0}$ and $(\mathcal{O}^{\times})^{r}$.
@@
We have $\phi(U_{S}\cap T_{1}^{0})\subset(\mathcal{O}^{1})^{r}$, for a proof see for example Lemma 7.3.16 in [19], where a much more general result is proven.
By Hensel’s Lemma, any element of $(\mathcal{O}^{1})^{r}$ has a square root inside $(\mathcal{O}^{1})^{r}$. Thus, we can find $t\in(\mathcal{O}^{1})^{r}$ such that $\phi(g^{-1}\sigma(g))=t^{2}$. We claim that $t\sigma(t)=1$. Denote $t^{\prime}=t\sigma(t)$, we know that $(t^{\prime})^{2}=1$. It is enough to check that $t^{\prime}$ is 1 modulo any power of $\pi$. As $t^{\prime}\in\mathcal{O}^{1}$, it is $1$ modulo $\pi$. Assume it is $1$ modulo $\pi^{k}$. Write $t^{\prime}\equiv 1+\pi^{k}b\text{ (mod }\pi^{k+1})$. We have $1\equiv t^{\prime 2}\equiv 1+2b\pi^{k}\text{ (mod }\pi^{k+1})$, so we get that $t^{\prime}$ is $1$ modulo $\pi^{k+1}$. By induction we are done.
Let $g^{\prime}=\phi^{-1}(t)\in T_{1}^{0}$. We have $g^{-1}\sigma(g)=(g^{\prime})^{2}=g^{\prime}\sigma(g^{\prime})^{-1}$. Another way to write this is as $gg^{\prime}=\sigma(gg^{\prime})$. Thus $gg^{\prime}\in H$. We know that $g^{\prime}\in T^{0}$, so $g^{\prime}$ acts trivially on $\mathcal{A}_{1}$. Thus, $gg^{\prime}\mathcal{A}_{1}=g\mathcal{A}_{1}=\mathcal{A}_{2}$ and $gg^{\prime}$ acts on $\mathcal{C}$ as $g$ does which means it fixes it pointwise and we are done.
@@ 3.1 I orbits on $X$
To describe the $H$ orbits on $G/I$, we recall that $G/I$ can be identified with the set of $\Omega$ colored chambers in the affine building $\mathcal{B}_{G}$. We denote the set of $\Omega$ colored chambers by $C_{\Omega}(G)=\{(\mathcal{C},o)|\mathcal{C}\subset\mathcal{B}_{G},o\in\Omega\}$. The group $G$ acts on $C_{\Omega}(G)$ by $g(\mathcal{C},o)=(g\mathcal{C},\omega(g)o)$.
Let $\mathbf{A}_{\sigma}$ be the set of $\sigma$ stable apartments, $H$ acts on $\mathbf{A}_{\sigma}$ and by Theorem 2.36 this action has finitely many orbits. Let $\mathcal{A}$ be a representative of an orbit and let $C_{\mathcal{A}}$ be the set of all $\Omega$ colored chambers contained in an apartment that is in the $H$ orbit of $\mathcal{A}$. Clearly $C_{\mathcal{A}}$ does not depend on the chosen representative and is stable under the action of $H$. By Proposition 3.1, the sets of the form $C_{\mathcal{A}}$ are pairwise disjoint.
It is enough to describe the $H$ orbits on $C_{\mathcal{A}}$.
@@
[…]
@@
It is clear that the map is well defined and surjective, we only have to check that the map is injective. Assume we have $w_{1},w_{2}\in W_{aff}(T)$ such that we can find $h\in H$ with $hw_{1}\mathcal{C}=w_{2}\mathcal{C}$ and $\omega(hw_{1})=\omega(w_{2})$, we need to show that $w_{1}w_{2}^{-1}\in HT^{0}$.
We have $hw_{1}w_{2}^{-1}w_{2}\mathcal{C}=w_{2}\mathcal{C}$. By Proposition 3.1 applied to the chamber $w_{2}\mathcal{C}$ we know that there is $h^{\prime}\in H$ that fixes $w_{2}\mathcal{C}$ pointwise such that $hw_{1}w_{2}^{-1}\mathcal{A}=h^{\prime}\mathcal{A}$.
Therefore $h^{\prime-1}hw_{1}w_{2}^{-1}$ stabilizes $\mathcal{A}$ and fixes the chamber $w_{2}\mathcal{C}$ pointwise (pointwise, because $\omega(h^{\prime})=\omega(hw_{1}w_{2}^{-1})=1)$. Thus $h^{\prime-1}hw_{1}w_{2}^{-1}\in T^{0}$. This completes the proof. ∎
[…]
@@
Let $I$ be the Iwahori subgroup that fixes a chamber contained in $\mathcal{A}_{0}$, for any $i\neq 0$ choose $g_{i}$ such that $g_{i}\mathcal{A}_{0}=\mathcal{A}_{i}$ and set $g_{0}=1$. Let $T$ be the torus that corresponds to $\mathcal{A}_{0}$.
From Proposition 3.3 we deduce:
@@
[…]
Theorem 3.4 allows us to define an action of the extended affine Weyl group of $G$ on the $I$ orbits on $X$. The group acts independently on each $W^{H}_{aff}(g_{i}Tg_{i}^{-1})$ and on each one it acts via an isomorphism between the abstract group $W_{aff}$ and $W_{aff}(g_{i}Tg_{i}^{-1})$, and the action of $W_{aff}(g_{i}Tg_{i}^{-1})$ on its quotient. We do not have a canonical isomorphism between $W_{aff}$ and $W_{aff}(g_{i}Tg_{i}^{-1})$. We use a chosen isomorphism between $W_{aff}$ and $W_{aff}(T)$ (given by Theorem 2.5) and conjugation by $g_{i}$ to identify $W_{aff}(T)$ and $W_{aff}(g_{i}Tg_{i}^{-1})$.
We also give a more explicit description of this action. Denote by $\mathcal{C}_{0}$ the chamber fixed pointwise by $I$, we have $\mathcal{C}_{0}\subset\mathcal{A}_{0}$.
First, notice that Theorem 3.4 implies the decomposition $G=\cup_{i=0}^{k}HW_{aff}(g_{i}Tg_{i}^{-1})g_{i}I=\cup_{i=0}^{k}Hg_{i}W_{aff}(T)I$.
Definition 3.5.
Let $x\in G$ be a representative of a double coset $HxI$. Let $w\in W_{aff}\cong W_{aff}(T)$. There are $g_{i}$ and $w_{1}\in W_{aff}(T)$ such that $HxI=Hg_{i}w_{1}I$. We define the action as:
$HxI\times w=Hg_{i}w_{1}wI$
We need to check that this is well defined and does not depend on the choice of representatives.
Proposition 3.6.
  1. 1.
    The double coset $Hg_{i}w_{1}wI$ depends only on $HxI$ and $w$.
  2. 2.
[…]
  3. 3.
    The defined action does not depend on the choice of the representatives $g_{i}$.
Proof.
Let $g^{\prime}\in G$ be such that $g^{\prime}\mathcal{A}_{0}$ is a $\sigma$ stable apartment and let $w^{\prime}_{1}\in W_{aff}$ be such that $Hg_{i}w_{1}I=Hg^{\prime}w^{\prime}_{1}I$. In order to prove items 1 and 3 it is enough to show that $Hg_{i}w_{1}wI=Hg^{\prime}w^{\prime}_{1}wI$. Notice that both $g_{i}w_{1}$ and $g^{\prime}w^{\prime}_{1}$ send $\mathcal{A}_{0}$ to a $\sigma$ stable apartment. Thus we can replace $g_{i}w_{1}$ with $g_{i}$ and $g^{\prime}w^{\prime}_{1}$ with $g^{\prime}$. It is enough to prove that if $Hg_{i}I=Hg^{\prime}I$ then we also have $Hg_{i}wI=Hg^{\prime}wI$. We now omit the $i$ index and write $g=g_{i}$.
Let $\mathcal{C}$ denote the chamber fixed pointwise by $I$. Note that $\mathcal{C}\subset\mathcal{A}_{0}$.
Let $h\in H$ such that $hg^{\prime}I=gI$. Denote $\mathcal{A}=g\mathcal{A}_{0}$ and $\mathcal{A}^{\prime}=hg^{\prime}\mathcal{A}_{0}$, both are $\sigma$ stable apartments that contain $g\mathcal{C}$. By Proposition 3.1 applied to the chamber $g\mathcal{C}$ and the apartments $\mathcal{A}$ and $\mathcal{A}^{\prime}$ we can find $h^{\prime}\in H$ such that $h^{\prime}g\mathcal{C}=g\mathcal{C}$ and $h^{\prime}\mathcal{A}^{\prime}=\mathcal{A}$. We also have $h^{\prime}hg^{\prime}I=h^{\prime}gI=gI$.
By replacing $h$ with $h^{\prime}h$ we can assume that $g\mathcal{A}_{0}=hg^{\prime}\mathcal{A}_{0}$.
Consider $x=g^{-1}hg^{\prime}$, we have $x\in I$ and $x\mathcal{A}_{0}=\mathcal{A}_{0}$, therefore $x$ must fix $\mathcal{A}_{0}$ pointwise. In particular $x$ fixes $w\mathcal{C}$ pointwise, which implies $w^{-1}xw=w^{-1}g^{-1}hg^{\prime}w\in I$ or $hg^{\prime}wI=gwI$. In particular $Hg^{\prime}wI=HgwI$.
Item 2 follows immediately from item 1. ∎
@@
Notice that by the construction of the action it follows that for $x\in H\backslash G/I$ and $w\in W_{aff}$ we can find a $\Omega$ colored chamber $(\mathcal{C}_{x},o_{x})\in C_{\Omega}(G)$ whose $H$ orbit is $x$ and another $\Omega$ colored chamber $(\mathcal{C}_{x\times w},o_{x\times w})\in C_{\Omega}(G)$ whose $H$ orbits is $x\times w$ such that $\mathcal{C}_{x}$ and $\mathcal{C}_{x\times w}$ are contained in a single $\sigma$ stable apartment.